woolyfrog:

Use the table below to answer this question: x y 0 -1 2 2 1 6 Find the average rate of change for the given function from x = 0 to x = 2.

woolyfrog:

@Zepdrix

woolyfrog:

@dude @Pixel

dude:

\(\frac{subtract~the~y~values}{subtract~the~x~values}\)

woolyfrog:

1 for the x values and -9 for the y values?

Zepdrix:

at x=0, the corresponding y is -1, we can write this as the ordered pair (0,-1) and at x=2, we have the ordered pair (2, 2)

Zepdrix:

\[\large\rm \text{average rate of change}=\frac{y_2-y_1}{x_2-x_1}\]

Zepdrix:

\[\large\rm =\frac{2-(-1)}{2-0}\]

woolyfrog:

yes i see

woolyfrog:

1/2?

Zepdrix:

minus a negative, can be rewritten as addition 2-(-1) = 2 + 1

woolyfrog:

3/2

Zepdrix:

Yayyy good job

woolyfrog:

whats next

Zepdrix:

The average rate of change from x=0 to x=2 is 3/2, or simply 1.5.

Zepdrix:

That's all c:

woolyfrog:

these are my choices A. 2 B. 4 C. −4 D. 1

woolyfrog:

@Zepdrix

Zepdrix:

Hmmmm weird :U

woolyfrog:

so which should i choose

Zepdrix:

It's 3/2. Are you sure you pasted the chart correctly? :O Hmm

woolyfrog:

yes

woolyfrog:

your right its the wrong chart

woolyfrog:

x y 0 2 1 -1 2 -6

woolyfrog:

@Zepdrix

woolyfrog:

-4!